OpenStudy (anonymous):

@jim_thompson5910 ' . Determine the slope between the points (-1, 0) and (-2, 3) (Points : 2) -2 -1 3 -3 I PUT B AM I RIGHT?

jimthompson5910 (jim_thompson5910):

Use the slope formula m = (y2-y1)/(x2-x1) m = ( 3- 0)/(-2-(-1)) m = ( 3- 0)/(-2+1) I'll let you finish

jimthompson5910 (jim_thompson5910):

what did you get this time

OpenStudy (anonymous):

ya :)

jimthompson5910 (jim_thompson5910):

slope = ya ?

OpenStudy (anonymous):

i got -1

jimthompson5910 (jim_thompson5910):

no the answer isn't -1

OpenStudy (anonymous):

D:

jimthompson5910 (jim_thompson5910):

-2+1 = ???

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

that's just part of it though m = (y2-y1)/(x2-x1) m = ( 3- 0)/(-2-(-1)) m = ( 3- 0)/(-2+1) m = 3/(-1) m = ???

OpenStudy (anonymous):

3

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

-3 you mean right?

jimthompson5910 (jim_thompson5910):

3 divide by -1 = -3

jimthompson5910 (jim_thompson5910):

since positive/negative = negative

OpenStudy (anonymous):

OH!

OpenStudy (anonymous):

TY :)

jimthompson5910 (jim_thompson5910):

yeah silly mistake, but it's ok